OpenStudy (hungergames13):

What is the approximate volume of a sphere with a radius of 9 in.? Use 3.14 to approximate pi and express your answer in hundredths.

OpenStudy (jdoe0001):

\(\textit{volume of a sphere}=\cfrac{4\pi r^3}{3}\qquad \begin{cases} r=radius \end{cases}\)